OpenStudy (anonymous):

can someone help with this problem -12 - -8 + -2 + +3

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

how did you get a -3?

OpenStudy (anonymous):

it is actually -12+8-2+3 = -4-2+3 = -6+3 = -3

OpenStudy (anonymous):

i got it; thank you

OpenStudy (anonymous):

u welcome..